Please remember to round your answer to two decimal places.

Problem 11-03 Net Salvage Value

Allen Air Lines must liquidate some equipment that is being replaced. The equipment originally cost $20 million, of which 85% has been depreciated. The used equipment can be sold today for $6 million, and its tax rate is 30%. What is the equipment's after-tax net salvage value? Write out your answer completely. For example, 2 million should be entered as 2,000,000.

$________

Problem 11-06 New-Project Analysis

The Campbell Company is considering adding a robotic paint sprayer to its production line. The sprayer's base price is $1,010,000, and it would cost another $17,500 to install it. The machine falls into the MACRS 3-year class (the applicable MACRS depreciation rates are 33.33%, 44.45%, 14.81%, and 7.41%), and it would be sold after 3 years for $677,000. The machine would require an increase in net working capital (inventory) of $20,000. The sprayer would not change revenues, but it is expected to save the firm $378,000 per year in before-tax operating costs, mainly labor. Campbell's marginal tax rate is 30%.

  1. What is the Year 0 net cash flow? $__________
  2. What are the net operating cash flows in Years 1, 2, and 3? Do not round intermediate calculations. Round your answers to the nearest dollar.
    Year 1 $_____
    Year 2 $_____
    Year 3 $_____
  3. What is the additional Year 3 cash flow (i.e, the after-tax salvage and the return of working capital)? Do not round intermediate calculations. Round your answer to the nearest dollar. $__________
  4. If the project's cost of capital is 10 %, what is the NPV of the project? Do not round intermediate calculations. Round your answer to the nearest dollar.
    $ __________
    Should the machine be purchased?__________ (Yes/No)

Problem 12-01 AFN equation

Broussard Skateboard's sales are expected to increase by 20% from $8.0 million in 2016 to $9.60 million in 2017. Its assets totaled $5 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 7%, and the forecasted payout ratio is 60%. Use the AFN equation to forecast Broussard's additional funds needed for the coming year. Round your answer to the nearest dollar. Do not round intermediate calculations.

$__________

Problem 12-02 AFN equation

Broussard Skateboard's sales are expected to increase by 15% from $7.0 million in 2016 to $8.05 million in 2017. Its assets totaled $4 million at the end of 2016. Broussard is already at full capacity, so its assets must grow at the same rate as projected sales. At the end of 2016, current liabilities were $1.4 million, consisting of $450,000 of accounts payable, $500,000 of notes payable, and $450,000 of accruals. The after-tax profit margin is forecasted to be 7%, and the forecasted payout ratio is 60%. What would be the additional funds needed? Do not round intermediate calculations. Round your answer to the nearest dollar.
$_______________

Assume that an otherwise identical firm had $5 million in total assets at the end of 2016. Broussard's capital intensity ratio (A0*/S0) is __________ (higher than / lower than / equal to) than the otherwise identical firm; therefore, Broussard is __________ (less / more / the same) capital intensive - it would require __________ (a smaller / a larger / the same) increase in total assets to support the increase in sales.

Please remember to round your answer to two decimal places.
Answer #1

Answer:

Question 11.03)

given

Original cost=$20 million

depreciated =85%

Accumulated depreciation =85%*20=$17 million

Actual salvage value=20-17=$ 3 million

Resell rate=$ 6 million

Taxable income=6-3=$3 million

Tax=30%*3=$0.9 million

After tax salvage value =6-0.9=$5.1 million